# Dynamics of Pseudoentanglement

### v1

@@ Abstract
The dynamics of quantum entanglement plays a central role in explaining the emergence of thermal equilibrium in isolated many-body systems. However, entanglement is notoriously hard to measure, and can in fact be “forged”: recent works have introduced a notion of pseudoentanglement describing ensembles of many-body states that, while only weakly entangled, cannot be efficiently distinguished from states with much higher entanglement, such as random states in the Hilbert space. In this work we initiate the study of the dynamical generation and propagation of pseudoentanglement. As generic quantum dynamics tends to maximize actual entanglement, we consider constrained models of time evolution: automaton (i.e. reversible classical) circuits that, when fed suitable input states, provably produce the “standard models” of pseudoentangled ensembles—uniformly random subset-phase states and subset states—at late times, a phenomenon we name pseudothermalization. We examine (i) how a pseudoentangled (random subset-phase) ensemble on a small subsystem spreads to the whole system as a function of time, and (ii) how a pseudoentangled (random subset) ensemble is generated from an initial product state. We map the above problems onto a family of classical Markov chains on subsets of the computational basis. The mixing times of such Markov chains are related to the time scales at which the states produced from the dynamics become indistinguishable from Haar-random states at the level of each statistical moment (or number of copies). Based on a combination of rigorous bounds and conjectures supported by numerics, we argue that each Markov chain’s relaxation time and mixing time have different asymptotic behavior in the limit of large system size. This is a necessary condition for a cutoff phenomenon: an abrupt dynamical transition to equilibrium. We thus conjecture that our random circuits give rise to asymptotically sharp pseudothermalization transitions.
I Introduction.
@@
Figure 1: Main ideas of this work. (a) Quantum thermalization: chaotic quantum dynamics on an isolated many-body systems (modeled by a Haar-random unitary circuit, bottom) produces random states at late times, with near-maximal entanglement entropy described by the Page curve (top). (b) In analogy with thermalization, we consider families of random ‘automaton’ circuits (bottom, see text) that, acting on suitable ensembles of initial states, produce pseudoentangled state ensembles at late times. These have relatively low entanglement entropy (top), but are computationally indistinguishable from Page states. (c) We study the dynamical process by which pseudoentanglement is generated and/or propagated over time, quantified by the distinguishability of the output state ensembles from the Haar ensemble: can an observer, by making joint measurements on $m$ copies of an output state, efficiently tell it apart from a random state? For each value of $m$, this defines a time scale dubbed ($m$-copy) pseudothermalization time (vertical dashed lines). This time scale may become sharp in the thermodynamic limit if the problem features a cutoff phenomenon.
@@
Entanglement is a fundamental concept in quantum information science and in quantum many-body physics, being both a necessary resource for quantum advantage in information processing Horodecki and Horodecki (2009); Chitambar and Gour (2019) and a framework to characterize different phases of matter beyond conventional order parameters Kitaev and Preskill (2006); Li and Haldane (2008); Jiang et al. (2012). Furthermore, the generation and propagation of entanglement in many-body systems is a key ingredient in our understanding of how thermal equilibrium emerges from the otherwise reversible unitary dynamics of isolated systems Rigol et al. (2008); Nandkishore and Huse (2015); Kaufman et al. (2016); Mezei and van der Schee (2020). Due to entanglement, local subsystems may be described by mixed states even when the global state is pure. Such local mixed states can approach universal forms at late times in generic dynamics—namely the Gibbs ensemble $\rho_{A,{\rm th}}\propto{\rm Tr}_{\bar{A}}(e^{-\beta H})$ Alhambra (2022) and its generalizations Vidmar and Rigol (2016)—thus recovering predictions of thermodynamics. In the absence of any conservation laws (e.g. in driven systems without symmetries) the predicted late-time equilibrium state for each local subsystem is maximally mixed: $\rho_{A,{\rm th}}\propto\mathbb{I}_{A}$. This requires each local subsystem to become maximally entangled with the remainder of the system. More specifically, in systems modeled by random unitary circuits Nahum et al. (2017); Bertini et al. (2019); Zhou and Nahum (2020); Brandao et al. (2021); Fisher et al. (2022); Potter and Vasseur (2022); Mittal and Hunter-Jones (2023) (and even in some quasi-periodically driven systems Pilatowsky-Cameo et al. (2023, 2024)) one can show that the global pure state of the system resembles a typical Haar-random state in the Hilbert space, also known as a Page state, with a characteristic scaling of entanglement known as the Page curve Page (1993), sketched in Fig. 1(a).
@@
Despite its ubiquity and far-reaching importance, entanglement is generally difficult to control, quantify, and measure Horodecki and Horodecki (2009), especially beyond few-body settings Amico et al. (2008). The last decade has seen substantial progress in these directions, particularly thanks to experimental advances in quantum simulators and theoretical developments in randomized measurements Islam et al. (2015); Kaufman et al. (2016); Elben et al. (2018); Brydges et al. (2019); Huang et al. (2020); Elben et al. (2023); nonetheless, the cost of measuring the entanglement entropy of an $N$-qubit system remains exponential in $N$, ultimately limiting the reach of all these approaches.
@@
Recent works Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023) highlight the intrinsic subtlety of measuring entanglement in a dramatic way, by introducing a notion of pseudoentanglement—a forgery of entanglement that cannot be efficiently distinguished from the real thing. More specifically, they construct ensembles of quantum states that, while only weakly entangled, are indistinguishable from Page states to any observer with a polynomial amount of resources. These ensembles are interesting from the point of view of computation and cryptography Ji et al. (2018); Lu et al. (2023), and can be efficiently constructed through the use of quantum-secure pseudorandom permutations Aaronson et al. (2023); Zhandry (2016).
Motivated by the role of entanglement in quantum thermalization, in this work we consider the dynamical generation and propagation of pseudoentanglement in many-body systems. Clearly, to avert the generic tendency of interacting many-body systems to maximize entanglement, we must specialize to non-generic models of time evolution; however, within these constraints (to be discussed below), we introduce random circuit models that prepare pseudoentangled state ensembles as their equilibrium steady states, much like the Page states for generic quantum dynamics. This is sketched in Fig. 1(b). We name this dynamical approach to pseudoentangled ensembles pseudothermalization. Depending on one’s perspective, this may be viewed as either a spoofing of quantum thermalization that is practically indistinguishable from the genuine phenomenon, or as a perfectly valid instance of quantum thermalization based on an operational definition (i.e., if a deviation from thermal equilibrium cannot be efficiently detected, one may argue it does not count).
Specifically, we consider the random subset-phase state ensemble Aaronson et al. (2023) and the random subset state ensemble Giurgica-Tiron and Bouland (2023), which are fixed points of ‘automaton’ quantum circuits Gopalakrishnan (2018); Iaconis et al. (2019); Iaconis (2021); Klobas et al. (2021) (i.e., permutations of the computational basis, not to be confused with ‘quantum cellular automata’). We therefore introduce families of automaton circuits and study the convergence of certain non-equilibrium initial states to these equilibrium fixed points, i.e., the phenomenon of pseudothermalization. We focus on two scenarios modeled after paradigmatic problems in ‘true’ quantum thermalization: (i) the spreading of pseudoentanglement from a small subsystem placed in contact with a larger, disentangled system, and (ii) the generation of pseudoentanglement from an initial product state.
@@
The key technical step in both cases is a mapping of the quantum pseudothermalization problem onto the equilibration of certain classical Markov chains over subsets of the computational basis—with the size of subsets related to the number of state copies $m$ available to the observer. This mapping allows us to make contact with a rich body of mathematical work on the equilibration of Markov chains on finite spaces, including the so-called cutoff phenomenon Diaconis and Shahshahani (1981); Diaconis (1996); Chen and Saloff-Coste (2008); Lacoin and Leblond (2011); Chen (2021). This phenomenon refers to the sudden equilibration that can be seen in random walks over highly-connected graphs (for example permutations of a deck of cards induced by a riffle shuffle Aldous and Diaconis (1986)), in contrast with e.g. diffusion on a Euclidean lattice where equilibration is gradual and characterized by a dynamical scaling collapse $\sim t/L^{z}$, $z=2$.
@@
Based on this connection, we argue that pseudothermalization in our random circuit models happens in an asymptotically sharp fashion: our output states are easily distinguishable from Page states for a long time, until they abruptly (i.e., over a parametrically shorter window of time) become indistinguishable at the level of $m$ copies, Fig. 1(c). This dynamical transition in the distinguishability of low-entanglement states from Page states is a new fundamental phenomenon in the dynamics of isolated quantum many-body systems, that enriches and complements our growing understanding of thermal equilibrium at the level of higher statistical moments Roberts and Yoshida (2017); Hunter-Jones (2019); Cotler et al. (2023); Ippoliti and Ho (2022); Pilatowsky-Cameo et al. (2023); Fava et al. (2023) and opens a number of exciting directions for future work.
@@
The rest of the paper is organized as follows. In Sec. II we provide a brief, self-contained review of some useful background: pseudoentanglement (Sec. II.1), Markov chains on finite sets and the cutoff phenomenon (Sec. II.2), and certain random walks over permutation groups (Sec. II.3). We then introduce the general framework of pseudothermalization via automaton circuit dynamics in Sec. III, and apply it to the problem of spreading and generation of pseudoentanglement in Sec. IV and V respectively. Finally we summarize our results and list outstanding questions and future directions in Sec. VI.
II Review of key concepts
@@
In this work we make extensive use of the following asymptotic notation. Given two functions $f,g:\mathbb{N}\mapsto\mathbb{R}$, we write:
@@
II.1 Pseudoentanglement
An ensemble of quantum states on $N$ qubits is said to be pseudoentangled if it cannot be distinguished from the unitarily invariant (Haar) random ensemble by looking at any ${\rm poly}(N)$ number of copies Aaronson et al. (2023). More precisely, defining the $m$-th moment operator
@@ II.1 Pseudoentanglement
an ensemble $\mathcal{E}$ is pseudoentangled [^1] The definition in Ref. Aaronson et al. (2023) additionally requires the ability to prepare the ensemble efficiently through suitable pseudorandom primitives. We will not focus on that aspect in this work. if
$\left\|\rho_{\mathcal{E}}^{(m)}-\rho_{{\rm Haar}}^{(m)}\right\|_{\rm tr}=o\left(1/{\rm poly}(N)\right)$ (2)
whenever $m=O({\rm poly}(N))$. The trace norm $\|A\|_{\rm tr}={\rm Tr}(\sqrt{A^{\dagger}A})$ characterizes the distinguishability between two quantum states: if $\|\rho-\sigma\|_{\rm tr}=\epsilon$, then there is an observable222 One can write the observable as $\mathcal{O}=\Pi_{+}-\Pi_{-}$, where $\Pi_{\pm}$ are the projectors on positive (negative) eigenvalue subspaces of $\rho-\sigma$. $\mathcal{O}$, with $-I\leq\mathcal{O}\leq I$, that gives ${\rm Tr}(\mathcal{O}\rho)-{\rm Tr}(\mathcal{O}\sigma)=\epsilon$; then an observer can tell apart $\rho$ from $\sigma$ by measuring $\mathcal{O}$ a number of times $\sim 1/\epsilon^{2}$. This is the optimal way to distinguish two states Nielsen and Chuang (2010). Thus the meaning of Eq. (2) is that, when making joint measurements on $m$ copies of states drawn from $\mathcal{E}$, an observer would need a superpolynomial number of experiments to tell the two ensembles apart. Since we require Eq. (2) to apply for all $m\leq O({\rm poly}(N))$, the states from $\mathcal{E}$ necessarily appear to be Haar-random (and thus near-maximally entangled) to any observer with polynomial resources, even though they may in fact be much less entangled.
A simple pseudoentangled ensemble is given by random subset phase states Aaronson et al. (2023)
@@
where $f$ is a random Boolean function on $\{0,1\}^{N}$ chosen uniformly at random, and $S$ is a subset of bitstrings, $S\subset\{0,1\}^{N}$, chosen uniformly at random out of all subsets of some fixed cardinality $|S|=K$. We call this space of subsets $\Sigma_{K}$. Here $K$ must scale suitably with the system size $N$: $\omega({\rm poly}(N))\leq K\leq o(2^{N})$. The role of $K$ is to determine the actual entanglement of the states in the ensemble: we have the average von Neumann entropy
@@
for all subsystems $A$ (regardless of locality), due to the fact that $K$ upper bounds the Schmidt rank of |$\ket{\psi_{S,f}}$ on any bipartition of the $N$ qubits. Thus the restriction on $K\geq\omega({\rm poly}(N))$ ensures that, e.g., the half-cut entanglement entropy can scale faster than $\log N$ (necessary so that the states cannot be efficiently distinguished from Haar-random ones by measuring their purity), and the restriction $K\leq o(2^{N})$ ensures that it scales more slowly than the Page value $\simeq N\log(2)/2$ Page (1993) (see Fig. 1(a,b)).
Very recently, Ref. Giurgica-Tiron and Bouland (2023) showed that the random phase factors are not necessary to have a pseudoentangled ensemble. Random subset states suffice:
@@
$\|\rho_{\mathcal{E}}^{(m)}-\rho_{{\rm Haar}}^{(m)}\|_{{\rm tr}}\leq O\left(\frac{m^{2}}{K}\right)+O\left(\frac{mK}{D}\right),$ (6)
@@
II.2 Equilibration of Markov chains and cut-off phenomenon
Figure 2: Schematic of the cutoff phenomenon for Markov chains on finite sets. (a) A family of Markov chains on systems of size $N$ has a cutoff if it exhibits “sudden” equilibration as $N\to\infty$, in the following sense: the total variation distance from equilibrium remains close to maximal for a time scale $t_{\rm mix}$ (mixing time), then decays to zero exponentially with a time constant $t_{\rm rel}$ (relaxation time) with $\lim_{N\to\infty}t_{\rm rel}/t_{\rm mix}=0$. (b) Even though the relaxation time itself may be growing with $N$, upon rescaling time by $t\mapsto t/t_{\rm mix}$ the approach to equilibrium becomes step-like as $N\to\infty$. In this sense there is a dynamically sharp equilibration transition.
A key technical step in this work will be to map the evolution of quantum distinguishability measures such as Eq. (2) to the equilibration of classical Markov chains over finite sets. For this purpose it is useful to briefly review some notations and facts about this topic, including the cut-off phenomenon which will play an important role in our results.
A Markov chain on a finite set can be simply defined by a stochastic matrix $\Gamma_{ij}$, with $i,j$ ranging over the set. ‘Stochastic’ means that $\Gamma_{ij}\geq 0$, with $\sum_{i}\Gamma_{ij}=1$ for all $j$. $\Gamma_{ij}$ represents the probability of a transition from $j$ to $i$. A state of the Markov chain is given by a probability distribution $p(i)$. In each time step, $p$ is updated according to $p(i)\mapsto\sum_{j}\Gamma_{ij}p(j)$. We say that the Markov chain equilibrates if $\Gamma^{t}p$ converges to some late-time limit $\pi$ as $t\to\infty$.
The equilibration of finite-sized Markov chains is characterized by two fundamental time scales Levin and Peres (2017). The first is the mixing time,
$t_{\rm mix}(\epsilon)=\min\{t:\|p_{t}-\pi\|_{\rm tv}\leq\epsilon\},$ (7)
describing the approach to the steady distribution. Here $\|\cdots\|_{\rm tv}$ is the total variation norm, $\|p\|_{\rm tv}=\frac{1}{2}\sum_{i}|p(i)|$, and one should maximize over all possible initial states. The second important time scale is the relaxation time, defined as the inverse spectral gap of the transition matrix $\Gamma_{ij}$:

### v3

@@ Abstract
The dynamics of quantum entanglement plays a central role in explaining the emergence of thermal equilibrium in isolated many-body systems. However, entanglement is notoriously hard to measure, and can in fact be “forged”: recent works have introduced a notion of pseudoentanglement describing ensembles of many-body states that, while only weakly entangled, cannot be efficiently distinguished from states with much higher entanglement, such as random states in the Hilbert space. This prompts the question: how much entanglement is truly necessary to achieve thermal equilibrium in quantum systems? In this work we address this question by introducing random circuit models of quantum dynamics that, at late times, equilibrate to pseudoentangled ensembles—a phenomenon we name ensemble pseudothermalization. These models replicate all the efficiently observable predictions of thermal equilibrium, while generating only a small (and tunable) amount of entanglement, thus deviating from the “maximum-entropy principle” that underpins thermodynamics. We examine (i) how a pseudoentangled ensemble on a small subsystem spreads to the whole system as a function of time, and (ii) how a pseudoentangled ensemble can be generated from an initial product state. We map the above problems onto a family of classical Markov chains on subsets of the computational basis. The mixing times of such Markov chains are related to the time scales at which the states produced from the dynamics become indistinguishable from Haar-random states at the level of each statistical moment, or number of copies. Based on a combination of rigorous bounds and conjectures supported by numerics, we argue that each Markov chain’s relaxation time and mixing time have different asymptotic behavior in the limit of large system size. This is a necessary condition for a cutoff phenomenon: an abrupt dynamical transition to equilibrium. We thus conjecture that our random circuits give rise to asymptotically sharp distinguishability transitions.
arxiv: 2403.09619
1 Introduction.
@@
Figure 1: Main ideas of this work. (a) Quantum thermalization: chaotic quantum dynamics on an isolated many-body systems (modeled by a Haar-random unitary circuit, bottom) produces random states at late times, with near-maximal entanglement entropy described by the Page curve (top). (b) In analogy with thermalization, we consider families of random ‘automaton’ circuits (bottom, see text) that, acting on suitable ensembles of initial states, produce pseudoentangled state ensembles at late times. These have relatively low entanglement entropy (top), but are computationally indistinguishable from Page states. (c) We study the dynamical process by which pseudoentanglement is generated and/or propagated over time, quantified by the distinguishability of the output state ensembles from the Haar ensemble: can an observer, by making joint measurements on $m$ copies of an output state, efficiently tell it apart from a random state? For each value of $m$, this defines a time scale dubbed $m$-copy indistinguishability time (vertical dashed lines). This time scale may become sharp in the thermodynamic limit if the problem features a cutoff phenomenon.
@@
Entanglement is a fundamental concept in quantum information science and in quantum many-body physics: it is at once a resource for quantum advantage in information processing horodecki_quantum_2009 ; chitambar_quantum_2019 and a framework to characterize different phases of matter beyond conventional order parameters kitaev_topological_2006 ; li_entanglement_2008 ; jiang_identifying_2012 . Furthermore, the generation and propagation of entanglement in many-body systems is a key ingredient in our understanding of how thermal equilibrium emerges from the otherwise reversible unitary dynamics of isolated systems rigol_thermalization_2008 ; nandkishore_many-body_2015 ; kaufman_quantum_2016 ; mezei_black_2020 . Due to entanglement, local subsystems may be described by mixed states even when the global state is pure; these local mixed states can equilibrate to universal forms alhambra_quantum_2022 ; vidmar_generalized_2016 , thus recovering the predictions of thermodynamics. In the absence of any conservation laws (for example in driven systems without symmetries), the predicted equilibrium is simply the maximally-mixed state $\rho_{A,{\rm th}}\propto\mathbb{I}_{A}$ in each local subsystem $A$, indicating saturation of entanglement to its maximal value. Informally, at late times the state of the system is expected to resemble a typical Haar-random state in the Hilbert space [^1] Convergence to Haar-random states can be shown rigorously in systems modeled by random unitary circuits nahum_quantum_2017 ; bertini_entanglement_2019 ; zhou_entanglement_2020 ; brandao_models_2021 ; fisher_random_2023 ; potter_entanglement_2022 ; mittal_local_2023 , and even in some quasi-periodically driven systems pilatowsky-cameo_complete_2023 ; pilatowsky-cameo_hilbert-space_2024 ., with a characteristic extensive scaling of entanglement known as the Page curve Page_average_1993 , sketched in Fig. 1(a).
@@
Despite its ubiquity and far-reaching importance, entanglement is generally difficult to control, quantify, and measure horodecki_quantum_2009 , especially beyond few-body settings amico_entanglement_2008 . The last decade has seen substantial progress in these directions, particularly thanks to experimental advances in quantum simulators and theoretical developments in randomized measurements islam_measuring_2015 ; kaufman_quantum_2016 ; elben_renyi_2018 ; brydges_probing_2019 ; huang_predicting_2020 ; elben_randomized_2023 ; nonetheless, the cost of measuring the entanglement entropy of an $N$-qubit system remains exponential in $N$ in the worst case, ultimately limiting the reach of all these approaches.
@@
Recent works aaronson_quantum_2023 ; giurgica-tiron_pseudorandomness_2023 ; jeronimo_pseudorandom_2024 highlight the intrinsic subtlety of measuring many-body entanglement in a dramatic way, by introducing a notion of pseudoentanglement—a forgery of entanglement that cannot be efficiently distinguished from the real thing. More specifically, these works construct ensembles of quantum states that, while only weakly entangled, are indistinguishable from random states (whose entropy tracks the Page curve and is thus near-maximal) to any observer with a polynomial amount of resources (time, memory, etc). These ensembles are interesting from the point of view of computation and cryptography ji_pseudorandom_2018 ; lu_quantum_2023 , and can be efficiently constructed through the use of quantum-secure pseudorandom permutations aaronson_quantum_2023 ; zhandry_note_2016 .
Motivated by these new developments, in this work we reexamine the role of entanglement in quantum thermalization.
1.1 Summary of results
We introduce constrained models of time evolution whose late-time equilibrium is not described by the highly-entangled Page states usually associated to high-temperature thermalization, but rather by pseudoentangled state ensembles. These ensembles are weakly entangled, yet they reproduce all the efficiently verifiable predictions of thermal equilibrium. We term this condition (and the dynamical processes that generate it over time) “ensemble pseudothermalization”. Depending on one’s point of view, this may be viewed as a well-disguised imitation of thermal equilibrium, or as a perfectly valid instance of it. The former perspective can be argued on first principles: pseudoentangled state ensembles by definition do not obey the “maximum entropy principle” that underlies thermodynamics [^2] More precisely, they do obey a maximum entropy principle, but only within a restricted space characterized by low entanglement. and are not stable against generic perturbations. The latter perspective can be argued on operational grounds: if a deviation from thermal equilibrium cannot be detected efficiently, it should not count. It is also worth noting that this type of equilibration happens at the level of an ensemble, for example over realizations of a random circuit, rather than at the level of individual quantum states as in conventional quantum thermalization srednicki_chaos_1994 ; rigol_thermalization_2008 ; nandkishore_many-body_2015 . Pseudoentanglement is an ensemble property; there is no sense in which a single quantum state can be pseudoentangled.
Understanding the process of ensemble pseudothermalization requires studying the dynamical generation and propagation of pseudoentanglement in many-body systems. This is analogous to the role of entanglement dynamics in thermalization (with the above caveat about the necessity of an ensemble). Like in that case, random circuit models of time evolution prove very useful nahum_quantum_2017 ; fisher_random_2023 ; potter_entanglement_2022 . We introduce random circuit models that prepare pseudoentangled state ensembles as their equilibrium steady states, much like the Page states for generic quantum dynamics. This is sketched in Fig. 1(b). Specifically, we consider the random subset-phase state ensemble aaronson_quantum_2023 and the random subset state ensemble giurgica-tiron_pseudorandomness_2023 ; jeronimo_pseudorandom_2024 . We introduce families of random reversible classical circuits (also known as ‘automaton’ quantum circuits [^3] Automaton quantum circuits should not be confused with ‘quantum cellular automata’, which are general locality-preserving unitaries farrelly_review_2020 . in the condensed matter literature gopalakrishnan_operator_2018 ; iaconis_anomalous_2019 ; iaconis_quantum_2021 ; klobas_exact_2021 ) that have pseudoentangled state ensembles as their late-time fixed points, and use them to quantitatively study the process of ensemble pseudothermalization.
@@ 1.1 Summary of results
We focus on two scenarios modeled after paradigmatic problems in ‘true’ quantum thermalization: (i) the spreading of pseudoentanglement from a small subsystem placed in contact with a larger, disentangled system, and (ii) the generation of pseudoentanglement from an initial product state. The key technical step in both cases is a mapping of pseudoentanglement generation onto the equilibration of certain classical Markov chains over subsets of the computational basis—with the size of subsets related to the number of state copies $m$ available to the observer. This mapping allows us to make contact with a rich body of mathematical work on the equilibration of Markov chains on finite spaces, including the so-called cutoff phenomenon diaconis_generating_1981 ; diaconis_cutoff_1996 ; chen_cutoff_2008 ; lacoin_cutoff_2011 ; chen_cutoff_2021 . This phenomenon refers to the sudden equilibration that can be seen in certain random walks, often over highly-connected graphs (for example permutations of a deck of cards induced by a riffle shuffle aldous_shuffling_1986 ), in contrast with e.g. diffusion on a Euclidean lattice where equilibration is smooth and characterized by a dynamical scaling collapse $\sim t/L^{z}$, $z=2$.
@@
Based on this connection, we argue that ensemble pseudothermalization in our random circuit models happens in an asymptotically sharp fashion: our output states are easily distinguishable from random states for a long time, until they abruptly (i.e., over a parametrically shorter window of time) become indistinguishable at the level of $m$ copies, Fig. 1(c). This dynamical transition in the distinguishability of low-entanglement states from random states is a new fundamental phenomenon in the dynamics of isolated quantum many-body systems. It enriches and complements our growing understanding of thermal equilibrium at the level of higher statistical moments roberts_chaos_2017 ; hunter-jones_unitary_2019 ; cotler_emergent_2023 ; ippoliti_solvable_2022 ; pilatowsky-cameo_complete_2023 ; fava_designs_2023 and opens a number of exciting directions for future work.
Our results imply several interesting consequences for the hardness of learning complex features of quantum dynamics, such as entanglement growth and measurement-induced phenomena, and pose limits on the existence of general, efficient protocols to address these problems.
1.2 Structure of the paper
@@
The rest of the paper is organized as follows. In Sec. 2 we provide a brief, self-contained review of some useful background: pseudoentanglement (Sec. 2.1), Markov chains on finite sets and the cutoff phenomenon (Sec. 2.2), and certain random walks over permutation groups (Sec. 2.3). We then introduce the general framework of ensemble pseudothermalization via automaton circuit dynamics in Sec. 3, and apply it to the problem of spreading and generation of pseudoentanglement in Sec. 4 and 5 respectively. Finally we summarize our results and discuss their practical and fundamental implications, as well as outstanding questions and future directions, in Sec. 6.
2 Review of key concepts
@@
In this work we make extensive use of the following asymptotic notation. Given two functions $f,g:\mathbb{N}\mapsto\mathbb{R}^{+}$, we write:
@@
2.1 Pseudoentanglement
An ensemble $\mathcal{E}$ of quantum states on $N$ qubits is said to be pseudoentangled aaronson_quantum_2023 if the states |$\ket{\psi}\in\mathcal{E}$ meet the following criteria [^4] Ref. aaronson_quantum_2023 gives a more general definition based on two ensembles with different entanglement that are computationally indistinguishable from each other. Here we always take the second ensemble to be the Haar-random ensemble. (to be specified precisely below):
  1. (i)
    they are efficiently preparable;
  2. (ii)
    they have low entanglement;
  3. (iii)
    they are computationally pseudorandom.
[…]
@@ 2.1 Pseudoentanglement
the ensemble $\mathcal{E}$ is computationally pseudorandom if, for any efficient quantum algorithm $\mathcal{A}$, we have
$\left|\mathcal{A}\left(\rho_{\mathcal{E}}^{(m)}\right)-\mathcal{A}\left(\rho_{{\rm Haar}}^{(m)}\right)\right|=o\left(1/{\rm poly}(N)\right)$ (2)
whenever $m=O({\rm poly}(N))$. In words, this means that any polynomial-time observation on a polynomial number of state copies cannot distinguish $\mathcal{E}$ from the Haar ensemble by more than a superpolynomially small amount, $\epsilon=o(1/{\rm poly}(N))$. Statistically resolving this discrepancy between the two ensembles would require at least $\Omega(1/\epsilon^{2})=\omega({\rm poly}(N))$ repetitions of the experiment. Thus, with a polynomially bounded amount of time and memory, nothing the observer can do would successfully distinguish $\mathcal{E}$ from the Haar distribution.
In this work we will drop the restriction on the efficiency of the algorithm $\mathcal{A}$, allowing arbitrary measurements on the $m$ copies. Then Eq. (2) reduces to
$D_{\rm tr}\left(\rho_{\mathcal{E}}^{(m)},\rho_{{\rm Haar}}^{(m)}\right)=o\left(1/{\rm poly}(N)\right),$ (3)
where $D_{\rm tr}(\rho,\sigma)=\frac{1}{2}\|\rho-\sigma\|_{\rm tr}$ is the trace distance and $\|A\|_{\rm tr}={\rm Tr}(\sqrt{A^{\dagger}A})$ is the trace norm. The trace distance characterizes the distinguishability between two quantum states: if $D_{\rm tr}(\rho,\sigma)=\epsilon$, then there is an observable666 One can write the observable as $\mathcal{O}=\Pi_{+}-\Pi_{-}$, where $\Pi_{\pm}$ are the projectors on positive (negative) eigenvalue subspaces of $\rho-\sigma$. $\mathcal{O}$, with $-I\leq\mathcal{O}\leq I$, that gives ${\rm Tr}(\mathcal{O}\rho)-{\rm Tr}(\mathcal{O}\sigma)=2\epsilon$. An observer can tell apart $\rho$ from $\sigma$ by measuring $\mathcal{O}$ a number of times $\sim 1/\epsilon^{2}$. This is the optimal way to distinguish two states nielsen_quantum_2020 . Eq. (3) describes “information-theoretic” pseudorandomness, a condition stronger than computational pseudorandomness. It describes an ensemble $\mathcal{E}$ that cannot be distinguished from the Haar ensemble by making a polynomial number of experiments on polynomially many copies, even if the experiments are allowed arbitrary complexity (e.g. projective measurements in a basis prepared by an exponentially deep circuit).
A simple pseudoentangled ensemble is given by random777 If $f$ and $S$ are taken to be genuinely random, then the ensemble satisfies Eq. (3) (information-theoretic pseudorandomness), but is not efficiently preparable. However one can replace them by suitable pseudorandom functions and permutations, respectively, in such a way that the ensemble is efficiently preparable and satisfies Eq. (2) (computational pseudorandomness). subset phase states aaronson_quantum_2023 ,
@@
where $f$ is a random Boolean function on $\{0,1\}^{N}$, and $S$ is a subset of bitstrings, $S\subset\{0,1\}^{N}$, chosen uniformly at random out of all subsets of some fixed cardinality $|S|=K$. We call this space of subsets $\Sigma_{K}$. To ensure pseudoentanglement, $K$ must scale suitably with the system size $N$: $\omega({\rm poly}(N))\leq K\leq o(2^{N/2})$. The role of $K$ is to determine the actual entanglement of the states in the ensemble: we have the average von Neumann entropy
@@
for all subsystems $A$ (regardless of locality), due to the fact that $K$ upper bounds the Schmidt rank of |$\ket{\psi_{S,f}}$ on any bipartition of the $N$ qubits. Thus the restriction on $K\geq\omega({\rm poly}(N))$ ensures that, e.g., the half-cut entanglement entropy can scale faster than $\log N$ (necessary so that the states cannot be efficiently distinguished from Haar-random ones by measuring their purity), and the restriction $K\leq 2^{o(N)}$ ensures sub-volume-law scaling [see Fig. 1(a,b)].
Very recently, Refs. giurgica-tiron_pseudorandomness_2023 ; jeronimo_pseudorandom_2024 showed that the random phase factors are not necessary to have a pseudoentangled ensemble. Random subset states suffice:
@@
$D_{\rm tr}\left(\rho_{\mathcal{E}}^{(m)},\rho_{{\rm Haar}}^{(m)}\right)\leq O\left(\frac{m^{2}}{K}\right)+O\left(\frac{mK}{D}\right),$ (7)
@@
2.2 Equilibration of Markov chains and cut-off phenomenon
Figure 2: Schematic of the cutoff phenomenon for Markov chains on finite sets. (a) A family of Markov chains on systems of size $N$ has a cutoff if it exhibits “sudden” equilibration as $N\to\infty$, in the following sense: the total variation distance from equilibrium remains close to maximal for a time scale $t_{\rm mix}$ (mixing time), then decays to zero exponentially with a time constant $t_{\rm rel}$ (relaxation time) with $\lim_{N\to\infty}t_{\rm rel}/t_{\rm mix}=0$. (b) Even though the relaxation time itself may be growing with $N$, upon rescaling time by $t\mapsto t/t_{\rm mix}$ the approach to equilibrium becomes step-like as $N\to\infty$. In this sense the dynamics has a sharp equilibration transition.
A key technical step in this work will be to map the evolution of quantum distinguishability measures, Eq. (3), to the equilibration of classical Markov chains over finite sets. For this purpose it is useful to briefly review some notations and facts about this subject.
A Markov chain on a finite set can be simply defined by a stochastic matrix $\Gamma_{ij}$, with $i,j$ ranging over the set. ‘Stochastic’ means that $\Gamma_{ij}\geq 0$ for all $i,j$, with $\sum_{i}\Gamma_{ij}=1$ for all $j$. $\Gamma_{ij}$ represents the probability of a transition from $j$ to $i$. A state of the Markov chain is given by a probability distribution $p(i)$. In each time step, $p$ is updated according to $p(i)\mapsto\sum_{j}\Gamma_{ij}p(j)$. We say that the Markov chain equilibrates if $\Gamma^{t}p$ converges to a late-time steady state $\pi$ as $t\to\infty$ (in this work we will always deal with Markov chains that admit a unique steady state).
The equilibration of finite-sized Markov chains is characterized by two fundamental time scales levin_markov_2017 . The first is the mixing time,
$t_{\rm mix}(\epsilon)=\min\{t:\,D_{\rm tv}(p_{t},\pi)\leq\epsilon\},$ (8)
describing the approach to the steady distribution. Here $D_{\rm tv}$ is the total variation distance, $D_{\rm tv}(p,q)=\frac{1}{2}\sum_{i}|p(i)-q(i)|$, and one should maximize over all possible initial states. The second important time scale is the relaxation time, defined as the inverse spectral gap of the transition matrix $\Gamma_{ij}$: